# Robust Conformal Outlier Detection under Contaminated Reference Data

### v1

@@ Abstract
Keywords: Conformal Prediction, Hypothesis Testing, Out-of-Distribution Detection, Contaminated Data
@@ 1.1 Background and Motivation
The broad need for reliable ML systems has sparked a surge of interest in conformal prediction—a versatile framework that can provide statistical guarantees for any “black-box” predictive model [38]. This framework formulates the outlier detection task as a statistical test, where the null hypothesis is that the new data point is not an outlier [23, 5]. To derive a decision rule guaranteeing type-I error control, conformal inference relies on a reference (calibration) set of inlier data points. These points are assumed to be sampled i.i.d. from an unknown distribution, independent of the data used to train the outlier detection model.
In practice, however, it is often difficult to obtain a perfectly clean reference dataset that contains no outliers [28, 44, 8, 20]. In our example, a more realistic scenario would assume instead access to a slightly contaminated reference set, mostly legitimate transactions with a few unnoticed outliers [44]. But this setting poses new challenges for conformal prediction methods, potentially invalidating the error control guarantees or, as we shall see, often reducing the power to detect true outliers at test time.
@@ 1.2 Outline and Contributions
A straightforward approach to cleaning the contaminated set is to remove all data points flagged as likely outliers by the detection model. However, this method is unsatisfactory, as it risks inadvertently removing inliers along with outliers, resulting in an "overly clean" reference set. This, in turn, distorts the inlier distribution and inflates the type-I error rate above the desired nominal level.
@@
Finally, in Section 4, we empirically validate our theory and proposed data-cleaning approach through comprehensive experiments on real-world datasets. The experiments confirm that conformal inference with contaminated data tends to be conservative. Furthermore, they demonstrate that our method significantly boosts power, particularly when the target type-I error rate is low and the number of outliers in the contaminated set is small. Software for reproducing the experiments is available at https://github.com/Meshiba/robust-conformal-od.
@@ 1.3 Related Work
Recently, there has been growing interest in studying the statistical properties of conformal inference methods under more realistic scenarios, moving beyond the idealized assumption of perfectly clean and exchangeable observations to account for various types of distribution shift [36, 12, 33, 4, 17, 40, 16, 18, 30, 34, 31]. This paper draws inspiration from several prior works in this area.
Tibshirani et al. [36] introduced a weighted conformal prediction approach to address covariate shift between calibration and test data, later extended by Podkopaev and Ramdas [30] to accommodate label shift. Both settings, however, involve a different form of distribution shift from the one we study here. Barber et al. [4] extend this line of work by analyzing the effects of general distribution shifts on the validity of conformal methods, focusing however on worst-case scenarios; see also Farinhas et al. [14].
@@
A more closely related line of work investigates the robustness of conformal prediction to label noise [12, 33, 10, 29] or other forms of data contamination [41, 42, 15]. Specifically, Einbinder et al. [12] and Sesia et al. [33] show that, under certain assumptions, conformal prediction for classification with noisy labels often results in conservative type-I error rates. Furthermore, Sesia et al. [33] propose a method to address this conservativeness by leveraging an explicit “label noise model” that captures the relationship between the true and contaminated labels in the calibration dataset.
In contrast, this paper avoids relying on an explicit model for the contaminated data, as such models can be difficult to estimate in practice within our context. Instead, we utilize a pre-trained black-box outlier detection model and a limited annotation budget to selectively and reliably trim outliers from the contaminated set. Furthermore, it is important to emphasize that the method proposed by Sesia et al. [33] is primarily designed for classification tasks with relatively balanced data, whereas outlier detection naturally involves extreme class imbalance. This distinction underscores the need for solutions specifically tailored to outlier detection.
@@ 2.1 Inference with Clean Calibration Data
Proposition 2.1 (from Vovk et al. [38]).
@@ 2.3 Explaining the Conservativeness
We begin by examining Figure 1, which provides some insight into this behavior based on the analysis of the “shuttle” dataset [3], as detailed in Section 4. In this example, the nonconformity scores of outlier data points in the contaminated calibration set are typically larger than those of the inliers. This pattern reflects the goal of a well-designed outlier detection model: to differentiate outliers from inliers by assigning higher scores to the former. Consequently, conformal p-values computed using (2) will be inflated relative to the ideal scenario in which all calibration points are inliers, reducing our power to detect true outliers at test time.
Figure 1: Histogram of nonconformity scores for inliers and outliers in a contaminated calibration subset of the “shuttle" data, with a contamination rate of 5%. The vertical lines indicate the $(1-\alpha)$ empirical quantile of all calibration scores (black), as well as separately for inliers (blue) and outliers (red), with $\alpha=0.02$.
This phenomenon is further corroborated by extensive numerical experiments presented in Section 4 and Sections B.2 and B.3, which consistently demonstrate this conservative behavior across nine different datasets.
While the conservativeness of conformal prediction methods in the presence of contaminated calibration data has already been observed and studied theoretically in different contexts [12, 33, 10], prior works did not focus on outlier detection. Therefore, it is helpful to introduce a new theoretical result that precisely quantifies the inflation of standard conformal p-values that we often observe in practice. This will serve as a foundation for the novel method of computing adaptive conformal p-values presented in the next section.
@@
This result is related to Theorem 1 in Sesia et al. [33], which studies the behavior of conformal prediction sets for multi-class classification [24, 32] calibrated with contaminated data. The key distinction is in our treatment of the calibration set: we assume that $n_{0}$ and $n_{1}$ are fixed, whereas Sesia et al. [33] consider a mixture model where the observed proportions of data points from different classes in the calibration set are random. While treating $n_{0}$ and $n_{1}$ as fixed is convenient for this paper, we also include an additional result (Corollary A.3) in Appendix A, which reaches qualitatively similar conclusions by adopting an approach more closely aligned with Theorem 1 from Sesia et al. [33].
@@ 3.1 Key Idea: Boosting Power by Cleaning the Data
A large calibration set is often needed because the smallest conformal p-value obtainable through (2) scales as $1/n$. Thus, a large $n$ is critical for achieving high confidence in identifying outliers, especially in “needle-in-a-haystack” scenarios [5], where a few outliers must be detected in a large test set dominated by inliers. In such cases, the ability to obtain very small p-values is essential to achieve non-trivial power while controlling the false discovery rate [6].
@@ 3.2 A Simple but Unsatisfactory Approach: Naive-Trim
To address this challenge, we will now present a more sophisticated method, which we refer to as Label-Trim. These approach utilize a limited labeling budget to remove outliers from $\operatorname{\mathcal{D}}_{\mathrm{cal}}$ in a more reliable manner, mitigating the risk of over-correcting the conformal p-value.
@@ 3.3 The Label-Trim Method
Algorithms 1 and 2 summarize this procedure, which intuitively offers advantages over both the standard method for computing $\hat{p}_{n+1}$ in (2), by potentially increasing power, and the Naive-Trim approach, by mitigating the risk of over-correcting $\hat{p}_{n+1}$.
@@
$\operatorname{\mathbb{P}}\left(\hat{p}_{n+1}^{\mathrm{LT}}\leq\alpha\right)\leq\alpha+\frac{1}{n_{0}+1}-\operatorname{\mathbb{E}}\left[\frac{\hat{n}^{\mathrm{LT}}_{1}}{n_{0}+1}\left((1-\alpha)-\hat{F}_{1}^{\mathrm{LT}}\left(\hat{Q}_{1-\alpha}^{\mathrm{LT}}\right)\right)\right].$
@@ Methods
In all experiments, we randomly split a given dataset into disjoint training $\mathcal{D}_{\text{train}}$, calibration $\mathcal{D}_{\text{cal}}$, and test sets of inliers $\mathcal{D}_{\text{test}}^{\text{inlier}}$ and outliers $\mathcal{D}_{\text{test}}^{\text{outlier}}$. To simulate a realistic setting, we construct the training and contaminated calibration sets with the same contamination rate of $r\%$. The inlier $\mathcal{D}_{\text{test}}^{\text{inlier}}$ and outlier $\mathcal{D}_{\text{test}}^{\text{outlier}}$ test tests are used to compute the type-I error and power of the outlier detection model, respectively. To ensure fair comparisons, all conformal methods use the same outlier detection model, trained on $\mathcal{D}_{\text{train}}$. Performance metrics are evaluated across 100 random splits of the data. The size of each dataset, along with the details of how $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{cal}}$, and $\mathcal{D}_{\text{test}}$ are constructed are provided in Section B.1.
@@ 4.1 Tabular Data
We now compare the performance of the different methods on three benchmark tabular datasets for outlier detection, previously used in the conformal literature [5]. Since conclusions are similar across datasets, we focus here on results for the shuttle dataset [3]. Results for the credit card [2] and KDDCup99 [1] datasets are presented in Section B.2. For all conformal methods, we use Isolation Forest [25] as the base outlier detection model, implemented using scikit-learn with default hyperparameters [7].
Figure 2: Comparison of conformal outlier detection methods on a tabular dataset (“shuttle”) as a function of the contamination rate $r$ The target type-I error rate is $\alpha=0.02$. Left: Empirical type-I error. Middle: Average detection rate (power), where higher values indicate better performance. Right: Number of outliers trimmed by the Label-Trim method. Results are averaged across 100 random splits of the data.
Figure 2 presents the performance metrics of each method as a function of the contamination rate $r$. Following the left panel in that figure, we can see that the Standard conformal method results in conservative type-I error control, with a decrease in the error rate as the outlier proportion increases—a behavior that is aligned with Lemma 2.2. Notably, the type-I error of the Oracle method is tightly centered around $\alpha$, as guaranteed by Proposition 2.1. The Naive-Trim method does not control the type-I error rate, emphasizing the need for reliable data-cleaning procedures. In striking contrast, our Label-Trim method, achieves a valid type-I error rate. At lower outlier proportions, the empirical type-I error is close to $\alpha$, but the method becomes more conservative as the outlier proportion increases. This observation aligns with the upper bound on the error rate derived in Theorem 3.1. Notably, as the contamination rate in the training data increases, the outlier detection model’s ability to distinguish between inliers and outliers weakens. This, in turn, adversely affects the effectiveness of forming a subset of data points for annotation, as demonstrated in the right panel of Figure 2. The Small-Clean method also controls the type-I error but is more conservative than Label-Trim due to its much smaller reference set, which becomes even smaller as the contamination rate increases. Observe how the power of the Small-Clean method is lower than that of the Standard approach, despite the latter using a contaminated reference set. By contrast, our proposed Label-Trim method significantly improves the power of the Standard method and even achieves near-oracle performance when the outlier proportion is low.
@@ 4.2 Visual Data
In what follows, we compare all methods using benchmark visual datasets for outlier detection. Similar to Zhang et al. [43], we construct six datasets, where the inlier samples are always images from CIFAR10 [21, 22] and the outlier samples vary across datasets. Specifically, the outliers are drawn from (1) MNIST [11], (2) SVHN [27], (3) Texture [9], (4) Places365 [9], (5) TinyImageNet [37], and (6) CIFAR100 [22]. For all datasets, we use the outlier detection model proposed by Sun et al. [35], which operates on feature representations extracted by a pre-trained ResNet18 model. More details are in Section B.1.
Table 1(b) summarizes the results for all six datasets. Overall, we can see a trend similar to the one of the tabular data: the Standard and Small-Clean methods are valid but conservative, the Naive-Trim fails to control the type-I error, and our Label-Trim achieves a significant boost in power while practically controlling the type-I error. Notably, our Label-Trim method attains near-oracle performance for low contamination rates. Detailed results for each dataset are provided in Section B.3.

### v2

@@ Abstract
Machine Learning, ICML
@@ 1.1 Background and Motivation
The broad need for reliable ML systems has sparked a surge of interest in conformal prediction—a versatile framework that can provide statistical guarantees for any “black-box” predictive model (Vovk et al., 2005). This framework formulates the outlier detection task as a statistical test, where the null hypothesis is that the new data point is not an outlier (Laxhammar & Falkman, 2015; Bates et al., 2023). To derive a decision rule guaranteeing type-I error control, conformal inference relies on a reference (calibration) set of inlier data points. These points are assumed to be sampled i.i.d. from an unknown distribution, independent of the data used to train the outlier detection model.
In practice, however, it is often difficult to obtain a perfectly clean reference dataset that contains no outliers (Park et al., 2021; Zhao et al., 2019; Chalapathy & Chawla, 2019; Jiang et al., 2022). In our example, a more realistic scenario would assume instead access to a slightly contaminated reference set, mostly legitimate transactions with a few unnoticed outliers (Zhao et al., 2019). But this setting poses new challenges for conformal prediction methods, potentially invalidating the error control guarantees or, as we shall see, often reducing the power to detect true outliers at test time.
@@ 1.2 Outline and Contributions
A straightforward approach to cleaning the contaminated set is to remove all data points flagged as likely outliers by the detection model. However, this method is unsatisfactory, as it risks inadvertently removing inliers along with outliers, resulting in an ”overly clean” reference set. This, in turn, distorts the inlier distribution and inflates the type-I error rate above the desired nominal level.
@@
Finally, in Section 4, we empirically validate our theory and proposed data-cleaning approach through comprehensive experiments on real-world datasets. The experiments confirm that conformal inference with contaminated data tends to be conservative. Furthermore, they demonstrate that our method significantly boosts power, particularly when the target type-I error rate is low and the number of outliers in the contaminated set is small.
A software that implements the proposed method is available at https://github.com/Meshiba/robust-conformal-od.
@@ 1.3 Related Work
Recently, there has been growing interest in studying the statistical properties of conformal inference methods under more realistic scenarios, moving beyond the idealized assumption of perfectly clean and exchangeable observations to account for various types of distribution shift (Tibshirani et al., 2019; Einbinder et al., 2024; Sesia et al., 2024; Barber et al., 2023; Gibbs & Candès, 2021; Zaffran et al., 2022; Feldman et al., 2023; Gibbs & Candès, 2024; Podkopaev & Ramdas, 2021; Si et al., 2023; Prinster et al., ). This paper draws inspiration from several prior works in this area.
Tibshirani et al. (2019) introduced a weighted conformal prediction approach to address covariate shift between calibration and test data, later extended by Podkopaev & Ramdas (2021) to accommodate label shift. Both settings, however, involve a different form of distribution shift from the one we study here. Barber et al. (2023) extend this line of work by analyzing the effects of general distribution shifts on the validity of conformal methods, focusing however on worst-case scenarios; see also Farinhas et al. (2024).
@@
A more closely related line of work investigates the robustness of conformal prediction to label noise (Einbinder et al., 2024; Sesia et al., 2024; Clarkson et al., 2024; Penso et al., 2025) or other forms of data contamination (Zaffran et al., 2023, 2024; Feldman & Romano, 2024). Specifically, Einbinder et al. (2024) and Sesia et al. (2024) show that, under certain assumptions, conformal prediction for classification with noisy labels often results in conservative type-I error rates. Furthermore, Sesia et al. (2024) proposes a method to address this conservativeness by leveraging an explicit “label noise model” that captures the relationship between the true and contaminated labels in the calibration dataset.
In contrast, this paper avoids relying on an explicit model for the contaminated data, as such models can be difficult to estimate in practice within our context. Instead, we utilize a pre-trained black-box outlier detection model and a limited annotation budget to selectively and reliably trim outliers from the contaminated set. Furthermore, it is important to emphasize that the method proposed by Sesia et al. (2024) is primarily designed for classification tasks with relatively balanced data, whereas outlier detection naturally involves extreme class imbalance. This distinction underscores the need for solutions specifically tailored to outlier detection.
@@ 2.1 Inference with Clean Calibration Data
Proposition 2.1 (from Vovk et al. (2005)).
@@ 2.3 Explaining the Conservativeness
We begin by examining Figure 1, which provides some insight into this behavior based on the analysis of the “shuttle” dataset (shu, ), as detailed in Section 4. In this example, the nonconformity scores of outlier data points in the contaminated calibration set are typically larger than those of the inliers. This pattern reflects the goal of a well-designed outlier detection model: to differentiate outliers from inliers by assigning higher scores to the former. Consequently, conformal p-values computed using (2) will be inflated relative to the ideal scenario in which all calibration points are inliers, reducing our power to detect true outliers at test time.
Figure 1: Histogram of nonconformity scores for inliers and outliers in a contaminated calibration subset of the “shuttle” data, with a contamination rate of 5%. The vertical lines indicate the $(1-\alpha)$ empirical quantile of all calibration scores (black), as well as separately for inliers (blue) and outliers (red), with $\alpha=0.02$.
This phenomenon is further corroborated by extensive numerical experiments presented in Section 4 and Section B.2, which consistently demonstrate this conservative behavior across nine different datasets.
While the conservativeness of conformal prediction methods in the presence of contaminated calibration data has already been observed and studied theoretically in different contexts (Einbinder et al., 2024; Sesia et al., 2024; Clarkson et al., 2024), prior works did not focus on outlier detection. Therefore, it is helpful to introduce a new theoretical result that precisely quantifies the inflation of standard conformal p-values that we often observe in practice. This will serve as a foundation for the novel method of computing adaptive conformal p-values presented in the next section.
@@
This result is related to Theorem 1 in Sesia et al. (2024), which studies the behavior of conformal prediction sets for multi-class classification (Lei et al., 2013; Romano et al., 2020) calibrated with contaminated data. The key distinction is in our treatment of the calibration set: we assume that $n_{0}$ and $n_{1}$ are fixed, whereas Sesia et al. (2024) consider a mixture model where the observed proportions of data points from different classes in the calibration set are random. While treating $n_{0}$ and $n_{1}$ as fixed is convenient for this paper, we also include an additional result (Corollary A.3) in Appendix A, which reaches qualitatively similar conclusions by adopting an approach more closely aligned with Theorem 1 from Sesia et al. (2024).
@@ 3.1 Key Idea: Boosting Power by Cleaning the Data
A large calibration set is often needed because the smallest conformal p-value obtainable through (2) scales as $1/n$. Thus, a large $n$ is critical for achieving high confidence in identifying outliers, especially in “needle-in-a-haystack” scenarios (Bates et al., 2023), where a few outliers must be detected in a large test set dominated by inliers. In such cases, the ability to obtain very small p-values is essential to achieve non-trivial power while controlling the false discovery rate (Benjamini & Hochberg, 1995).
@@ 3.2 A Simple but Unsatisfactory Approach: Naive-Trim
To address this challenge, we will now present a more sophisticated method, which we refer to as Label-Trim. This approach utilizes a limited labeling budget to remove outliers from $\operatorname{\mathcal{D}}_{\mathrm{cal}}$ in a more reliable manner, mitigating the risk of over-correcting the conformal p-value.
@@ 3.3 The Label-Trim Method
Algorithms 1 and 2 summarize this procedure, which intuitively offers advantages over both the standard method for computing $\hat{p}_{n+1}$ in (2), by potentially increasing power, and the Naive-Trim approach, by mitigating the risk of over-correcting $\hat{p}_{n+1}$. A schematic illustration of the construction of $\operatorname{\mathcal{D}}_{\mathrm{cal}}^{\mathrm{LT}}$ by Algorithm 1 is shown in Figure 2.
Figure 2: A schematic illustration of the proposed active data-cleaning of the contaminated reference set (Algorithm 1). The approach begins by computing nonconformity scores of the contaminated reference data $\mathcal{D}_{\mathrm{cal}}$ using a pretrained outlier detection model, where blue circles denote inliers and red circles denote outliers. The scores are then sorted in increasing order, and the top $m$ samples—those most likely to be outliers—are selected for annotation. After removing the annotated outliers, the resulting (partially) cleaned set $\mathcal{D}_{\mathrm{cal}}^{\mathrm{LT}}$ is used for calibration.
@@
$\operatorname{\mathbb{P}}\left(\hat{p}_{n+1}^{\mathrm{LT}}\leq\alpha\right)\leq\alpha+\frac{1}{n_{0}+1}$
$\qquad-\operatorname{\mathbb{E}}\left[\frac{\hat{n}^{\mathrm{LT}}_{1}}{n_{0}+1}\left((1-\alpha)-\hat{F}_{1}^{\mathrm{LT}}\left(\hat{Q}_{1-\alpha}^{\mathrm{LT}}\right)\right)\right].$
@@
Figure 3: Comparison of conformal outlier detection methods on a tabular dataset (“shuttle”) as a function of the contamination rate $r$. The target type-I error rate is $\alpha=0.02$. Left: Empirical type-I error. Middle: Average detection rate (power), where higher values indicate better performance. Right: Number of outliers trimmed by the Label-Trim method. Results are averaged across 100 random splits of the data.
@@ Methods
In all experiments, we randomly split a given dataset into disjoint training $\mathcal{D}_{\text{train}}$, calibration $\mathcal{D}_{\text{cal}}$, and test sets of inliers $\mathcal{D}_{\text{test}}^{\text{inlier}}$ and outliers $\mathcal{D}_{\text{test}}^{\text{outlier}}$. To simulate a realistic setting, we construct the training and contaminated calibration sets with the same contamination rate of $r\%$. The inlier $\mathcal{D}_{\text{test}}^{\text{inlier}}$ and outlier $\mathcal{D}_{\text{test}}^{\text{outlier}}$ test sets are used to compute the type-I error and power of the outlier detection model, respectively. To ensure fair comparisons, all conformal methods use the same outlier detection model, trained on $\mathcal{D}_{\text{train}}$. Performance metrics are evaluated across 100 random splits of the data. The size of each dataset, along with the details of how $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{cal}}$, and $\mathcal{D}_{\text{test}}$ are constructed are provided in Section B.1.
@@ 4.1 Tabular Data
We now compare the performance of the different methods on three benchmark tabular datasets for outlier detection, previously used in the conformal literature (Bates et al., 2023). Since conclusions are similar across datasets, we focus here on results for the shuttle dataset (shu, ). Results for the credit card (cre, ) and KDDCup99 (KDD, ) datasets are presented in Section B.2.1. For all conformal methods, we use Isolation Forest (Liu et al., 2008) as the base outlier detection model, implemented using scikit-learn with default hyperparameters (Buitinck et al., 2013).
Figure 3 presents the performance metrics of each method as a function of the contamination rate $r$. Following the left panel in that figure, we can see that the Standard conformal method results in conservative type-I error control, with a decrease in the error rate as the outlier proportion increases—a behavior that is aligned with Lemma 2.2. Notably, the type-I error of the Oracle method is tightly centered around $\alpha$, as guaranteed by Proposition 2.1. The Naive-Trim method does not control the type-I error rate, emphasizing the need for reliable data-cleaning procedures. In striking contrast, our Label-Trim method achieves a valid type-I error rate. At lower outlier proportions, the empirical type-I error is close to $\alpha$, but the method becomes more conservative as the outlier proportion increases. This observation aligns with the upper bound on the error rate derived in Theorem 3.1. Notably, as the contamination rate in the training data increases, the outlier detection model’s ability to distinguish between inliers and outliers weakens. This, in turn, adversely affects the effectiveness of forming a subset of data points for annotation, as demonstrated in the right panel of Figure 3. The Small-Clean method also controls the type-I error but is more conservative than Label-Trim due to its much smaller reference set, which becomes even smaller as the contamination rate increases. Observe how the power of the Small-Clean method is lower than that of the Standard approach, despite the latter using a contaminated reference set. By contrast, our proposed Label-Trim method significantly improves the power of the Standard method and even achieves near-oracle performance when the outlier proportion is low.
@@
Additional experiments
In Section B.2.1, we extend the experiments presented above. These include evaluations under higher contamination levels and additional outlier detection models, such as One-Class SVM (Schölkopf et al., 2001) and Local Outlier Factor (Breunig et al., 2000). We also examine the effect of different outlier injection strategies and the robustness of the proposed method under test-time distribution shifts in the outlier population.
@@ 4.2 Visual Data
In what follows, we compare all methods using benchmark visual datasets for outlier detection. Similar to Zhang et al. (2024), we construct six datasets, where the inlier samples are always images from CIFAR10 (Krizhevsky & Hinton, 2009; Krizhevsky et al., 2009) and the outlier samples vary across datasets. Specifically, the outliers are drawn from (1) MNIST (Deng, 2012), (2) SVHN (Netzer et al., 2011), (3) Texture (Cimpoi et al., 2014), (4) Places365 (Cimpoi et al., 2014), (5) TinyImageNet (Torralba et al., 2008), and (6) CIFAR100 (Krizhevsky et al., 2009). For all datasets, we use the outlier detection model proposed by Sun et al. (2021), ReAct, which operates on feature representations extracted by a pre-trained ResNet-18 model. More details are in Section B.1.
Table 1(b) summarizes the results for all six datasets. Overall, we can see a trend similar to that of the tabular data: the Standard and Small-Clean methods are valid but conservative, the Naive-Trim fails to control the type-I error, and our Label-Trim achieves a significant boost in power while practically controlling the type-I error. Notably, our Label-Trim method attains near-oracle performance for low contamination rates. Detailed results for each dataset, along with experiments with additional outlier detection models—specifically, the ReAct method (Sun et al., 2021) with a pre-trained VGG-19 model and the SCALE method (Xu et al., 2024) with a ResNet-18 model—are provided in Section B.2.2.